# Fundamental multiplication formulas

### v1

@@ Abstract.
By embedding the Hecke algebra $\check{H}_{\boldsymbol{q}}$ of type $D$ into the Hecke algebra $H_{{\boldsymbol{q}},1}$ of type $B$ with unequal parameters $({\boldsymbol{q}},1)$, the ${\boldsymbol{q}}$-Schur algebras $S^{\kappa}_{\boldsymbol{q}}(n,r)$ of type $D$ is naturally defined as the endomorphism algebra of the tensor space with the $\check{H}_{\boldsymbol{q}}$-action restricted from the $H_{{\boldsymbol{q}},1}$-action that defines the $({\boldsymbol{q}},1)$-Schur algebra $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ of type $B$. We investigate the algebras $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ and $S^{\kappa}_{\boldsymbol{q}}(n,r)$ both algebraically and geometrically and describe their standard bases, dimension formulas and weight idempotents. Most importantly, we use the geometrically derived two sets of the fundamental multiplication formulas in $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ to derive multi-sets (9 sets in total!) of the fundamental multiplication formulas in $S^{\kappa}_{\boldsymbol{q}}(n,r)$.
@@
Supported partially by NSF grant DMS1801915.
@@ 1. Introduction
The $q$-Schur algebras of type $A$ were originated from two sources. One is in the quantum group theory and the other is in the representation theory of finite general linear groups. In [17], Jimbo used $q$-Schur algebras to describe the (quantum) Schur–Weyl duality, while in [9] Dipper–James used them to investigate representations of finite general linear groups in non-defining characteristic. Thus, $q$-Schur algebras play a central role in connecting representations of quantum linear groups with those of finite general linear groups.
@@
Although one naturally expected thirty years ago the existence of such a connection for other finite classical groups, the Schur–Weyl–Brauer duality and its quantum version for classical types other than $A$ cannot be used to establish such a connection. This is because the Brauer algebras and their quantum analogs, the so-called BMW algebras (see [7, Th. 10.2.5] and the references therein), have a very different representation theroy from those of Hecke algebras of type $B/C/D$. Thus, what we expected here is a possible existence of a kind of “Schur–Weyl–Hecke duality” for other classical types. This puzzle has finally been solved thirty years later!
@@
In their study of quantum symmetric pairs and their canonical basis theory, Bao–Wang [2] discovered a new Schur–Weyl dualiy which is called here the Schur–Weyl–Hecke duality. The Schur–Weyl–Hecke duality of type $B/C$ involves the twin $i$-quantum groups ${\bf U}^{\jmath}(n)$ and ${\bf U}^{\imath}(n)$ and the Hecke algebras of types $B$ and $C$. These $i$-quantum groups appear in the quantum symmetric pairs of type AIII with no black nodes in their Satake diagrams. The $q$-Schur algebras involved have a lot of applications; see, e.g., [4, 12, 18, 20, 21, 23]. Note that, for the type $D$ case, a commuting action of the quantum symmetric pairs ${\bf U}^{\jmath}(n)$, ${\bf U}^{\imath}(n)$ and Hecke algebras of type $D$ on tensor spaces was first observed by Ehrig and Stroppel [13] in their study of maximal parabolic category ${\mathcal{O}}$/isotropic grassmannians of $\mathfrak{so}_{2n}$. See also Bao’s work [1] It is natural to ask if there is a Schur–Weyl–Hecke duality in type $D$?
@@
In [14], Z. Fan and the second-named author investigated a natural geometric setting of the $q$-Schur algebras of type $D$ (see [19, Appendix] for the associated algebraic setting) and their limiting algebra $\mathcal{K}$, following the work [4]. Moreover, certain $i$-quantum groups are also introduced so that the type $D$ $q$-Schur algebras are their homomorphic images. However, it has been noticed that multiplication formulas in the type $D$ $q$-Schur algebras discussed in [14, §4] are incomplete (see [14, Prop. 4.3.2]). In other words, more cases need to be considered in the determination of the standard basis when the algebra is defined by $q$-permutation modules of the type $D$ Hecke algebra (see [19, Appendix]). Hence, the dimension formulas need to be revised and multiplication formulas about these new basis elements should be computed as well.
@@
The aim of the paper is to fill these gaps in [14] and [19] and to set up a foundation for a possible solution to the aformentioned question. It is interesting to note that, on the way to achieve the goal, we discovered more. The strategy here is new as we develop two parallel theories: a type $B$ theory in unequal parameters and a type $D$ theory. We then use a splitting process via either weight idempotents or geometric splitting to link the two.
@@
We now describe it in more details. We start with the algebraic approach developed in [11], where the Hecke algebra $\check{H}_{\boldsymbol{q}}$ of type $D_{r}$ is embedded in the Hecke algebra $H_{{\boldsymbol{q}},1}$ of type $B_{r}$ in unequal parameters. Interestingly, there is a geometric construction for $H_{q,1}$ using complete isotropic flag variety over the even-order orthogonal group ${\text{\rm O}}_{2r}(q)$. Then, we may present the $(q,1)$-Schur algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ of type $B$ both algebraically, using $q$-permutation modules, and geometrically, using pairs of isotropic partial flag variety over ${\text{\rm O}}_{2r}(q)$ and their associated convolution algebra (see Theorem 3.10). In this way, we may define the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ of type $D$ by restriction from $H_{{\boldsymbol{q}},1}$ to $\check{H}$ and identify it with the algebra $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined in [19] (see Proposition 4.2). Next, using a criterion for splitting an ${\text{\rm O}}_{2r}(q)$-orbit into two ${\text{\rm SO}}_{2r}(q)$-orbits, a new parametrisation of ${\text{\rm SO}}_{2r}(q)$-orbits is obtained (see (4.8.3)). We then introduce the standard basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$, indexed by (labelled) double cosets in the type $D$ Weyl group, and the defining orbital basis for the convolution algebra over ${\text{\rm SO}}_{2r}(q)$, indexed by ${\text{\rm SO}}_{2r}(q)$-orbits. Theorem 4.15 gives a geometric setting for $S_{q}^{\textsc{d}}(n,r)$ using a basis correspondence. This precise basis correspondence (or identification) in loc. cit. is the key to finding the fundamental multiplication formulas in these generic $q$-Schur algebras of type $D$ (Theorems 6.26.4 and 7.17.3) via a splitting process of the counterpart formulas of type $B$ given in Theorem 5.2.
@@
More precisely, we summarise the fundamental multiplication formulas in the paper as follows. (See notational list below.)
@@
Let $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ and $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ be the ${\boldsymbol{q}}$-Schur algebras of type $B$ and $D$, respectively, with their respective bases $\{e_{A}\mid\Xi\}$ and $\{\phi_{\check{\mathbb{A}}}\mid{\check{\mathbb{A}}}\in\check{\Xi}\}$. Suppose $B,C\in\Xi$ such that $B-E_{h,h+1}^{\theta}$ and $C-E_{h+1,h}^{\theta}$ $(h\in[1,n])$ are diagonal and, in $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$,
$e_{B}e_{A}=\delta_{\text{\rm co}(B),\text{\rm ro}(A)}\sum_{1\leq p\leq N}g_{h,A,p}e_{{}_{h}\!A_{p}},\quad e_{C}e_{A}=\delta_{\text{\rm co}(C),\text{\rm ro}(A)}\sum\limits_{1\leqslant p\leqslant N}g^{\prime}_{{}_{h,A,\overline{p}}}e_{{}_{h}\!A_{\overline{p}}}.$
(See (3.7.1) for the notation ${}_{h}\!A_{p},{}_{h}\!A_{\overline{p}}$.) If ${\check{\mathbb{A}}}$, ${\check{\mathbb{B}}}$, and $\check{\mathbb{C}}$ in $\check{\Xi}$ are obtained by the splitting process from $A,B,$ and $C$, respectively, then the structure constants for $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ appearing in $\phi_{{\check{\mathbb{B}}}}\phi_{\check{\mathbb{A}}}$ (resp., $\phi_{\check{\mathbb{C}}}\phi_{\check{\mathbb{A}}}$) are either some $g_{h,A,p}$ (resp. $g^{\prime}_{{}_{h,A,\overline{p}}}$ or $\frac{1}{2}g^{\prime}_{{}_{h,A,\overline{p}}}$).
We organise the paper as follows. In Section 2, we first introduce, following [10, (3.3.2)(2)] and [11, Rem. 2.6(1)], the algebraic definition for the type $D$ $q$-Schur algebras $S^{\kappa}_{{\boldsymbol{q}}}(n,r)$ associated with (or covering) the type $B$ algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ in unequal parameters. We then reinterpret $S^{\kappa}_{{\boldsymbol{q}}}(n,r)$ as the algebra $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined by Heck algebras of type $D$ and their $q$-permutation modules (see [19, (A.3.3)]). In Section 3, we present the geometric definition of $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$, largely following [4, 14]. In Section 4, we develop a precise labelling for those ${\text{\rm O}}_{2r}(q)$-orbits which sprit into two ${\text{\rm SO}}_{2r}(q)$-orbits. This allows us to label the standard basis for $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ both algebraically in terms of double cosets and geometrically in terms of ${\text{\rm SO}}_{2r}(q)$-orbits. Finally, in Section 5, the multiplication formulas in the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ of type $B$ (with unequal parameters) will be derived geometrically via convolution products as in [4]. In the last Sections 6 and 7, fundamental multiplication formulas for the ${\boldsymbol{q}}$-Schur algebra of type $D$ are derived using the weight idempotent and orbit splitting techniques. This includes especially those missing ones in [14] (see Remarks 4.16 and 7.4).
@@
    Index set of orbits of parabolic subgroups
@@
    Hecke algebras;
@@
    ${\boldsymbol{q}}$-Schur algebra in unequal/equal parameter;
@@
    Standard bases for ${\boldsymbol{q}}$-Schur algebras;
@@
    Index set for standard basis;
@@ 1.1. Acknowledgements
The second-named author is grateful to Chun-Ju Lai for pointing out a gap in the parametrization of SO-orbits in [14]; see [19, (A.4.2)], Remark 4.16, and (4.8.3).
The authors would like to thank the University of New South Wales and the University at Buffalo for their hospitality during the writing of the paper. The third-named author thanks the China Scholarship Council for financial support.
@@ 2. The ${\boldsymbol{q}}$-Schur algebras of type $D$: an algebraic setting
Let $W=W^{\textsc{b}}=W(B_{r})$ ($r\geq 2$) and ${\check{W}}=W^{\textsc{d}}=W(D_{r})$ ($r\geq 4$) be the Weyl groups of type $B_{r}$ and $D_{r}$ associated with the Dynkin diagrams:
@@
where $S=\{s_{1},\cdots,s_{r-1},s_{r}\}$, ${\check{S}}=\{s_{1},\cdots,s_{r-1},\varsigma_{r}\}$ with $\varsigma_{r}:=s_{r}s_{r-1}s_{r}$, and $\bar{S}=\{s_{1},\cdots,s_{r-1}\}$, which is isomorphic to the symmetric group ${\mathfrak{S}}_{r}$.
Let $t_{r}=s_{r}$, $t_{r-i}=s_{r-i}t_{r-i+1}s_{r-i}$, $u_{i}=t_{r}t_{r-i}$ for $11\leq i\leq r-$ 1, and let $C=\langle t_{1},\cdots,t_{r}\rangle$ and $\check{C}=\langle u_{1},\cdots,u_{r-1}\rangle$. Then, we have $W=C\rtimes\bar{W}$ and ${\check{W}}=\check{C}\rtimes\bar{W}$. Moreover, we may identify ${\bar{W}}$ with the symmetric group ${\mathfrak{S}}_{r}$ on $r$ letters.
@@
The inner automorphism ${\mathfrak{f}}:w\mapsto s_{r}ws_{r}$ on ${{W}}$ induces by restriction a graph automorphism $f$ on ${\check{W}}$ which flips the Coxeter graph:
@@
Let $H_{{\boldsymbol{q}},1}=H_{{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}}\otimes_{{\mathcal{A}}^{\prime}}\mathbb{Z}[{\boldsymbol{q}}]$ be the algebra obtained by base change from ${\mathcal{A}}^{\prime}$ to ${\mathcal{A}}:=\mathbb{Z}[{\boldsymbol{q}}]$ under specializing ${\boldsymbol{q}}^{\prime}$ to 1. Then $H_{{\boldsymbol{q}},1}$ is presented by $T_{1},\ldots,T_{r-1},T_{r}$ and relations
@@
The $\mathbb{Z}[{\boldsymbol{q}}]$-algebra $H_{{\boldsymbol{q}},1}$ contains a subalgebra $\check{H}:=H^{\textsc{d}}_{\boldsymbol{q}}$ which is isomorphic to the Hecke algebra of type $D_{r}$. Here $\check{H}$ is generated by $T_{1},\ldots,T_{r-1},T_{\varsigma}:=T_{r}T_{r-1}T_{r}$ with defining relations (1)–(3) in (2.0.4) together with:
@@
We will use the same notation as above for the bases $\{T_{w}\mid w\in W\}$ for $H_{{\boldsymbol{q}},1}$ and $\{T_{w}\mid w\in{\check{W}}\}$ for $\check{H}$.
@@
We use the following indexing set to label parabolic subgroups of $W$:
@@
The double subscripts in the notation $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$ indicate this unequal parameter setting. It is different from the algebra ${\mathcal{S}}^{\jmath}(n,r)$ with equal-parameter investigated in [2] and [12].

### v2

# The $q$-Schur algebras in type $D$, I: Fundamental multiplication formulas

@@ Abstract.
By embedding the Hecke algebra $\check{H}_{\boldsymbol{q}}$ of type $D$ into the Hecke algebra $H_{{\boldsymbol{q}},1}$ of type $B$ with unequal parameters $({\boldsymbol{q}},1)$, the ${\boldsymbol{q}}$-Schur algebras $S^{\kappa}_{\boldsymbol{q}}(n,r)$ of type $D$ is naturally defined as the endomorphism algebra of the tensor space with the $\check{H}_{\boldsymbol{q}}$-action restricted from the $H_{{\boldsymbol{q}},1}$-action that defines the $({\boldsymbol{q}},1)$-Schur algebra $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ of type $B$. We investigate the algebras $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ and $S^{\kappa}_{\boldsymbol{q}}(n,r)$ both algebraically and geometrically and describe their natural bases, dimension formulas and weight idempotents. Most importantly, we use the geometrically derived two sets of the fundamental multiplication formulas in $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$ to derive multi-sets (9 sets in total!) of the fundamental multiplication formulas in $S^{\kappa}_{\boldsymbol{q}}(n,r)$.
@@
Supported partially by NSF grant DMS1801915 and Simons Travel Support for Mathematicians.
@@ 1. Introduction
The $q$-Schur algebras of type $A$ originated from two sources. One originates from quantum group theory and the other from representation theory of finite general linear groups. In [17], Jimbo used $q$-Schur algebras to describe the (quantum) Schur–Weyl duality, while in [9] Dipper–James used them to investigate representations of finite general linear groups in non-defining characteristic. Thus, $q$-Schur algebras play a central role in connecting representations of quantum linear groups with those of finite general linear groups.
@@
Although one naturally expected thirty years ago the existence of such a connection for other finite classical groups, the Schur–Weyl–Brauer duality and its quantum version for classical types other than $A$ cannot be used to establish such a connection. This is because the Brauer algebras and their quantum analogs, the so-called BMW algebras (see [7, Th. 10.2.5] and the references therein), have a very different representation theory from those of Hecke algebras of type $B/C/D$. Thus, we hypothesized the existence of a ‘Schur-Weyl-Hecke duality’ for other classical types. This puzzle has finally been solved after thirty years!
@@
In their study of quantum symmetric pairs and their canonical basis theory, Bao–Wang [2] discovered a new Schur–Weyl duality which is called here the Schur–Weyl–Hecke duality. The Schur–Weyl–Hecke duality of type $B/C$ involves the twin $i$-quantum groups ${\bf U}^{\jmath}(n)$ and ${\bf U}^{\imath}(n)$ and the Hecke algebras of types $B$ and $C$. These $i$-quantum groups appear in the quantum symmetric pairs of type AIII with no black nodes in their Satake diagrams. The $q$-Schur algebras involved have numerous applications; see, e.g., [4, 12, 18, 20, 21, 23]. Note that, for the type $D$ case, a commuting action of the $i$-quantum groups ${\bf U}^{\jmath}(n)$, ${\bf U}^{\imath}(n)$ and Hecke algebras of type $D$ on tensor spaces was first observed by Ehrig and Stroppel [13] in their study of maximal parabolic category ${\mathcal{O}}$/isotropic Grassmannians of $\mathfrak{so}_{2n}$. See also Bao’s work [1]. It is natural to ask if there is a Schur–Weyl–Hecke duality in type $D$?
@@
In [14], Z. Fan and the second-named author investigated a natural geometric setting of the $q$-Schur algebras of type $D$ (see [19, Appendix] for the associated algebraic setting) and their limiting algebra $\mathcal{K}$, following the work [4]. Moreover, certain $i$-quantum groups are also introduced so that the type $D$ $q$-Schur algebras are their homomorphic images. However, it has been noticed that multiplication formulas in the type $D$ $q$-Schur algebras discussed in [14, §4] are incomplete (see [14, Prop. 4.3.2]). In other words, more cases need to be considered in the determination of the natural basis when the algebra is defined by $q$-permutation modules of the type $D$ Hecke algebra (see [19, Appendix]). Hence, the dimension formulas need to be revised and multiplication formulas about these new basis elements should be computed as well.
@@
The aim of the paper is to fill these gaps in [14] and [19] and to set up a foundation for a possible solution to the aforementioned question. It is interesting to note that, on the way to achieve the goal, we discovered more. The strategy here is new as we develop two parallel theories: a type $B$ theory in unequal parameters and a type $D$ theory. We then use a splitting process via either weight idempotents or geometric splitting to link the two.
@@
We now describe it in more details. We start with the algebraic approach developed in [11], where the Hecke algebra $\check{H}_{\boldsymbol{q}}$ of type $D_{r}$ is embedded in the Hecke algebra $H_{{\boldsymbol{q}},1}$ of type $B_{r}$ in unequal parameters. Interestingly, there is a geometric construction for $H_{q,1}$, where $q$ is a prime power, using complete isotropic flag variety over the even-order orthogonal group ${\text{\rm O}}_{2r}(q)$. Then, we may present the $({\boldsymbol{q}},1)$-Schur algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ of type $B$ both algebraically, using $q$-permutation modules, and geometrically, using pairs of isotropic partial flag variety over ${\text{\rm O}}_{2r}(q)$ and their associated convolution algebra (see Theorem 3.10). In this way, we may define the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ of type $D$ by restriction from $H_{{\boldsymbol{q}},1}$ to $\check{H}_{\boldsymbol{q}}$ and identify it with the algebra $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined in [19] (see Proposition 4.2). Next, using a criterion for splitting an ${\text{\rm O}}_{2r}(q)$-orbit into two ${\text{\rm SO}}_{2r}(q)$-orbits, a new parametrization of ${\text{\rm SO}}_{2r}(q)$-orbits is obtained (see (4.8.3)). We then introduce the natural basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$, indexed by (labelled) double cosets in the type $D$ Weyl group, and the defining orbital basis for the convolution algebra over ${\text{\rm SO}}_{2r}(q)$, indexed by ${\text{\rm SO}}_{2r}(q)$-orbits. Theorem 4.15 gives a geometric setting for $S_{q}^{\textsc{d}}(n,r)$ via a basis correspondence. This precise basis correspondence (or identification) in loc. cit. is the key to finding the fundamental multiplication formulas in these generic $q$-Schur algebras of type $D$ (Theorems 6.26.4 and 7.17.3) through a splitting process of the corresponding formulas of type $B$ given in Theorem 5.2.
@@
More precisely, we summarize the fundamental multiplication formulas in the paper as follows. (See notational list below.)
@@
Let $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ and $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ be the ${\boldsymbol{q}}$-Schur algebras of type $B$ and $D$, respectively, over the polynomial ring $\mathbb{Z}[{\boldsymbol{q}}]$ with their respective bases $\{e_{A}\mid A\in\Xi\}$ and $\{\phi_{\check{\mathbb{A}}}\mid{\check{\mathbb{A}}}\in\check{\Xi}\}$. Suppose $B,C\in\Xi$ such that $B-E_{h,h+1}^{\theta}$ and $C-E_{h+1,h}^{\theta}$ $(h\in[1,n])$ are diagonal and, in $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$,
$e_{B}e_{A}=\delta_{\text{\rm co}(B),\text{\rm ro}(A)}\sum_{1\leq p\leq N}g_{h,A,p}e_{{}_{h}\!A_{p}},\quad e_{C}e_{A}=\delta_{\text{\rm co}(C),\text{\rm ro}(A)}\sum\limits_{1\leq p\leq N}g^{\prime}_{{}_{h,A,\overline{p}}}e_{{}_{h}\!A_{\overline{p}}},$
for some $g_{h,A,p},g^{\prime}_{{}_{h,A,\overline{p}}}\in\mathbb{Z}[{\boldsymbol{q}}]$. (See (3.7.1) for the notation ${}_{h}\!A_{p},{}_{h}\!A_{\overline{p}}$.) If ${\check{\mathbb{A}}}$, ${\check{\mathbb{B}}}$, and $\check{\mathbb{C}}$ in $\check{\Xi}$ are obtained by a certain splitting process from $A,B,$ and $C$, respectively, then the structure constants for $S_{{\boldsymbol{q}}}^{\kappa}(n,r)$ appearing in $\phi_{{\check{\mathbb{B}}}}\phi_{\check{\mathbb{A}}}$ (resp., $\phi_{\check{\mathbb{C}}}\phi_{\check{\mathbb{A}}}$) are some of the $g_{h,A,p}$ (resp. $g^{\prime}_{{}_{h,A,\overline{p}}}$ or $\frac{1}{2}g^{\prime}_{{}_{h,A,\overline{p}}}$).
We organize the paper as follows. In Section 2, we first introduce, following [10, (3.3.2)(2)] and [11, Rem. 2.6(1)], the algebraic definition for the type $D$ $q$-Schur algebras $S^{\kappa}_{{\boldsymbol{q}}}(n,r)$ associated with (or covering) the type $B$ algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ in unequal parameters. We then reinterpret $S^{\kappa}_{{\boldsymbol{q}}}(n,r)$ as the algebra $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ defined by Heck algebras of type $D$ and their $q$-permutation modules (see [19, (A.3.3)]). In Section 3, we present the geometric definition of $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)$, largely following [4, 14]. In Section 4, we develop a precise labelling for those ${\text{\rm O}}_{2r}(q)$-orbits which split into two ${\text{\rm SO}}_{2r}(q)$-orbits. This allows us to label the natural basis for $S^{\textsc{d}}_{\boldsymbol{q}}(n,r)$ both algebraically in terms of double cosets and geometrically in terms of ${\text{\rm SO}}_{2r}(q)$-orbits. Finally, in Section 5, the multiplication formulas in the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}},1}^{\jmath}(n,r)$ of type $B$ (with unequal parameters) will be derived geometrically via convolution products as in [4]. In the last Sections 6 and 7, fundamental multiplication formulas for the ${\boldsymbol{q}}$-Schur algebra of type $D$ are derived using either weight idempotents or orbit splitting techniques. This includes especially those missing ones in [14] (see Remarks 4.16 and 7.4).
@@
    Index sets of orbits of parabolic subgroups
@@
    Hecke algebras in unequal/equal parameter;
@@
    ${\boldsymbol{q}}$-Schur algebras;
@@
    Natural bases for ${\boldsymbol{q}}$-Schur algebras;
@@
    Index sets for natural bases;
@@ 1.1. Acknowledgements
The second-named author is grateful to Chun-Ju Lai for pointing out a gap in the parametrization of ${\text{\rm SO}}_{2r}(q)$-orbits in [14]; see [19, (A.4.2)], Remark 4.16, and (4.8.3).
The authors express their gratitude to the University of New South Wales and the University at Buffalo for their hospitality during the preparation of this paper. Additionally, the third author acknowledges the financial support provided by the China Scholarship Council.
@@
We sincerely thank the referee for valuable comments and suggestions.
@@ 2. The ${\boldsymbol{q}}$-Schur algebras of type $D$: an algebraic setting
We start with the definitions of the Weyl groups, Hecke algebras, and ${\boldsymbol{q}}$-Schur algebres of type $B/D$. By identifying $W$ and its parabolic subroups with certain fixed-point subgroups of the symmetric group ${\mathfrak{S}}_{2r}$, we then index the natural (or double coset) basis for the ${\boldsymbol{q}}$-Schur algebra $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$ of type $B$ by a certain matrix set $\Xi_{N,2r}$. This matrix set will soon be used in Section 3 to label the ${\text{\rm O}}_{2r}(q)$-orbits in ${\mathcal{X}}\times{\mathcal{X}}$ for an $n$-step isotropic flag variety ${\mathcal{X}}$.
2.1. The Weyl groups $W/{\check{W}}$, Hecke algebras $H_{{\boldsymbol{q}},1}/\check{H}_{\boldsymbol{q}}$, and ${\boldsymbol{q}}$-Schur algebras $S^{\jmath}_{{\boldsymbol{q}},1}(n,r)/S^{\kappa}_{\boldsymbol{q}}(n,r)$
Let $W=W^{\textsc{b}}=W(B_{r})$ ($r\geq 2$) and ${\check{W}}=W^{\textsc{d}}=W(D_{r})$ ($r\geq 4$) denote the Weyl groups of type $B_{r}$ and $D_{r}$ associated with the Dynkin diagrams:
@@
where $S=\{s_{1},\cdots,s_{r-1},s_{r}\}$, ${\check{S}}=\{s_{1},\cdots,s_{r-1},\varsigma_{r}\}$ with $\varsigma_{r}:=s_{r}s_{r-1}s_{r}$, and $\bar{S}=\{s_{1},\cdots,s_{r-1}\}$. Note that $\bar{W}$ is isomorphic to the symmetric group ${\mathfrak{S}}_{r}$ on $r$ letters.
Let $t_{r}=s_{r}$, $t_{r-i}=s_{r-i}t_{r-i+1}s_{r-i}$, $u_{i}=t_{r}t_{r-i}$ for $11\leq i\leq r-$ 1, and let $C=\langle t_{1},\cdots,t_{r}\rangle$ and $\check{C}=\langle u_{1},\cdots,u_{r-1}\rangle$. Then, we have $W=C\rtimes\bar{W}$ and ${\check{W}}=\check{C}\rtimes\bar{W}$. Moreover, we often identify ${\bar{W}}$ with ${\mathfrak{S}}_{r}$ in the sequel.
@@
The inner automorphism ${\mathfrak{f}}:w\mapsto s_{r}ws_{r}$ on ${{W}}$ induces by restriction a graph automorphism ${\mathfrak{f}}$ on ${\check{W}}$ which flips the Coxeter graph:
@@
Let $H_{{\boldsymbol{q}},1}=H_{{\boldsymbol{q}},{\boldsymbol{q}}^{\prime}}\otimes_{{\mathcal{A}}^{\prime}}\mathbb{Z}[{\boldsymbol{q}}]$ be the algebra obtained by base change from ${\mathcal{A}}^{\prime}$ to ${\mathcal{A}}:=\mathbb{Z}[{\boldsymbol{q}}]$ under the specialization of ${\boldsymbol{q}}^{\prime}=1$. Then $H_{{\boldsymbol{q}},1}$ is presented by $T_{1},\ldots,T_{r-1},T_{r}$ and relations
@@
The $\mathbb{Z}[{\boldsymbol{q}}]$-algebra $H_{{\boldsymbol{q}},1}$ contains a subalgebra $\check{H}_{\boldsymbol{q}}:=H^{\textsc{d}}_{\boldsymbol{q}}$ which is isomorphic to the Hecke algebra of type $D_{r}$. Here $\check{H}_{\boldsymbol{q}}$ is generated by $T_{1},\ldots,T_{r-1},T_{\varsigma}:=T_{r}T_{r-1}T_{r}$ with defining relations (1)–(3) in (2.0.4) together with:
@@
We will use the same notation as above describing basis $\{T_{w}\mid w\in W\}$ for $H_{{\boldsymbol{q}},1}$ and basis $\{T_{w}\mid w\in{\check{W}}\}$ for $\check{H}_{\boldsymbol{q}}$. We often write $H=H_{{\boldsymbol{q}},1}$ and $\check{H}=H_{\boldsymbol{q}}$ for notational simplicity (see §4.1).
@@
We use the following indexing set to label (standard) parabolic subgroups of $W$:
@@
The double subscripts in the notation $S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)$ indicate this unequal parameter setting. It is different from the equal-parameter algebra $\mathbf{S}^{\jmath}$ or ${\mathcal{S}}^{\jmath}_{\mathcal{Z}}(n,r)$ investigated in [4] or in [12].